# ACT-R: Adaptive Camera Trajectories for Single-View 3D Reconstruction

### v2

@@ Appendix D Calculation of Random Trajectory
Following the same set up as our adaptive trajectories, we apply uniform azimuth angle change of $18^{\circ}$ and divide the camera trajectory orbit into four segments. Instead of setting a constant elevation increment stepsize within each segment, we randomly sample an incrementation step with $[-5^{\circ},5^{\circ}]$ for each time of elevation change. To ensure the final trajectory forms a close loop, we apply the same mirroring and negating mechanism between segments.
Appendix E Alternative Feature Encoders
@@
In our primary experiments, we utilized VGG16 [4] as the feature extractor to quantify semantic differences between slice images and the input image. To evaluate the impact of feature encoder architecture, we conducted ablation studies with DINOv2 [70] as an alternative feature extractor. For the DINOv2 implementation, we extracted features from the final intermediate layer, resulting in a feature representation $\phi(\cdot)\in\mathbb{R}^{1536\times 16\times 16}$. These features were then used to construct semantic difference maps following the same methodology applied to VGG features.
@@
Qualitative and quantitative comparisons between static camera trajectory (SV3D), adaptive trajectory with DINOv2, and adaptive trajectory with VGG16 are presented in Figure 12 and Table 2, respectively. Each method was evaluated using both NeUS [2] and IM [5] reconstruction pipelines.
@@
Figure 12: Qualitative visual comparisons between 3D reconstruction with static camera trajectory (SV3D), adaptive camera trajectory with DINOv2 (DINO) and adaptive camera trajectory with VGG16 (Ours) on the GSO dataset. Please zoom in for a closer inspection. Meshes are reconstructed with both NeuS and LRM; the latter is noted with the subscript. It can be seen that adaptive camera trajectory provides additional geometric awareness in comparison to static camera trajectory, regardless.
@@
SV3D(u) [1] 4.93 2.79 18.4
[…]
$\text{Ours}_{\text{DINO}}$ 4.59 3.16 16.9
$\text{Ours}_{\text{DINO}_{\text{IM}}}$ 4.55 3.40 16.4
Ours 4.42 3.52 16.6
$\text{Ours}_{\text{IM}}$ 4.51 3.42 16.2
Table 2: Quantitative results of single-view 3D reconstruction on the full GSO dataset with static camera trajectory and adaptive trajectories with DINOv2 [70] and VGG16 [4] as feature extractors, respectively. Numbers in bold are ranked first.
The results demonstrate that both feature extractors contribute to developing more effective camera paths, yielding improvements in both qualitative and quantitative metrics compared to static trajectories. Notably, VGG16 outperforms DINOv2 in this specific application. We hypothesize that this performance difference stems from VGG16’s superior capability in processing low-level geometric features present in our grayscale slice images, whereas DINOv2 was primarily trained on full-color RGB images and may be less attuned to the specific characteristics of our data.
Appendix F Orbital Camera Trajectories
@@
Appendix G Complementary Role in Direct 3D supervision

### v3

@@ 10 Calculation of Random Trajectory
Following the same setup as our adaptive trajectories, we apply a uniform azimuth angle change of $18^{\circ}$ and divide the camera trajectory orbit into four segments. Instead of setting a constant elevation increment stepsize within each segment, we randomly sample an incrementation step with $5^{\circ},5^{\circ}]$ for each time of elevation change. To ensure the final trajectory forms a close loop, we apply the same mirroring and negating mechanism between segments.
11 Semantic Difference Map
We show additional examples on semantic difference map obatained by comparing feature differences between input image and the generated slices from Slice3D [64], as shown in Figure 11.
Figure 11: Visualization of semantic difference map computed form feature comparison between input image and the generated slice images by Slice3D [64]. The brighter color indicates higher semantic difference, and the darker colour indicates lower semantic differences.
12 Alternative Feature Encoders
@@
In our primary experiments, we utilized VGG16 [52] as the feature extractor to quantify semantic differences between slice images and the input image. To evaluate the impact of feature encoder architecture, we conducted ablation studies with DINOv2 [42] and CLIP [46] as alternative feature extractors. For the DINOv2 implementation, we extracted features from the final intermediate layer, resulting in a feature representation $\phi(\cdot)\in\mathbb{R}^{1536\times 16\times 16}$ while CLIP features remain in the same resolution as VGG. These features were then used to construct semantic difference maps following the same methodology applied to VGG features.
@@
Qualitative and quantitative comparisons between static camera trajectory (SV3D), adaptive trajectory with CLIP, DINOv2, and VGG are presented in Figure 12 and Table 2, respectively. All methods were evaluated on the IM [73] reconstruction pipeline.
@@
Figure 12: Qualitative visual comparisons between 3D reconstruction with static camera trajectory (SV3D), adaptive camera trajectory with CLIP [46], DINOv2 [42] (DINO) and VGG16 (Ours) as the feature extractor on the GSO dataset. Please zoom in for a closer inspection. All meshes are reconstructed with LRM. It can be seen that adaptive camera trajectory provides additional geometric awareness in comparison to static camera trajectory, regardless.
@@
$\text{SV3D}_{\text{IM}}$ [60] 4.60 7.74 18.0
$\text{Ours}_{\text{CLIP}_{\text{IM}}}$ 3.93 9.99 16.1
$\text{Ours}_{\text{DINO}_{\text{IM}}}$ 3.97 9.91 16.3
$\text{Ours}_{\text{IM}}$ 3.93 9.93 16.1
Table 2: Quantitative results of single-view 3D reconstruction on the full GSO dataset with static camera trajectory and adaptive trajectories with CLIP [46], DINOv2 [42] and VGG16 [52] as feature extractors, respectively. Numbers in bold are ranked first.
The results demonstrate that all feature extractors contribute to developing more effective camera paths, yielding improvements in both qualitative and quantitative metrics compared to static trajectories. Among the different feature encoder choices, VGG16 and CLIP exhibit similar performance with minimal differences in F1 scores, while DINOv2 falls slightly short.
13 Coverage Metric
We further evaluate static, random, and adaptive trajectories using a coverage metric. Given the ground truth (GT) mesh as input, we render the GT mesh at 21 camera poses for each trajectory type. The visible area in each render is marked as visible on the corresponding mesh texture UV map, and the final coverage metric is measured as the accumulated visible region over the entire texturable UV map area. Averaged across all objects in the GSO dataset, the static trajectory achieved an 87.8% coverage rate, random trajectory 89.0%, adaptive trajectory with CLIP 90.3%, VGG16 90.5%, and DINOv2 93.4%.
While the coverage metric indicates the occlusion-revelation capability of different trajectories to some extent, we emphasize that it involves a trade-off with reconstruction quality. For example, using a greedy algorithm for extreme occlusion-revelation path-finding can achieve near 100% coverage, but it induces drastic elevation changes and degrades video generation quality due to sudden viewpoint transitions. Considering both metrics, feature encoders including VGG16 and CLIP strike an effective balance between coverage and quality.
14 Orbital Camera Trajectories
@@
15 Complementary Role in Direct 3D supervision
@@
Figure 14: Additional qualitative visual comparisons between single-view 3D reconstruction methods on the GSO dataset. Please zoom in for a closer inspection. Meshes reconstructed with NeUS and LRM are in yellow and blue blocks, respectively. Our slice-guided approach demonstrates superior reconstruction of occluded regions, including bowl bottoms and extended features like elephant trunk. Please zoom in for a detailed comparison of the reconstruction quality.
Figure 15: Additional qualitative visual comparisons between single-view 3D reconstruction methods on the GSO dataset. Please zoom in for a closer inspection. Meshes reconstructed with NeUS and LRM are in yellow and blue blocks, respectively. Our slice-guided approach demonstrates superior reconstruction of occluded regions, including bowl bottoms, hidden wings and raccoon tails. Please zoom in for a detailed comparison of the reconstruction quality.